Please held me solve this problem:
Mark is 11 years older that his sister. In 8 years he will be twice as old as she will be then. How old are they now?

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

Please held me solve this problem:
Mark is 11 years older that his sister. In 8 years he will be twice as old as she will be then. How old are they now?
------------------
Ages NOW:
Sister's age is "x"
Mark's age is "x+11"
------------------------
Ages in 8 years:
Sister will be "x+8"
Mark will be "x+19"
------------------------
EQUATION:
x+19 = 2(x+8)
x+19 = 2x+16
x = 3
---------
Sister is NOW 3 years old
Mark is NOW 14 years old
===========================
Cheers,
Stan H.

Answer by checkley77(12844)

M=S+11
M+8=2(S+8)
(S+11)+8=2(S+8)
S+19=2S+16
S-2S=16-19
-S=-3
S=3 YEARS OLD FOR THE SISTER NOW.
M=3+11=14 YEARS OLD FOR MARK NOW.
PROOF:
14+8=2(3+8)
22=2*11
22=22